Compensation=$329
Consumption=$370
Export=$55
Net foriegn factor income =$3
Government spending=$43
Gross investment=$43
Import=$63
Net interest=$49
Profit=$69
Rental income=$46
Calculate GDP and GNP with both approaches to national income accounting?
1
Expert's answer
2020-06-24T13:41:07-0400

Expenditure Approach

The gross domestic product is computed using the following formula:

"\text{GDP=Consumption+Government spending+Investment+(Export - Imports)}"

"\text{GDP}=370+43+43+(55-63)"

"\text{GDP}=448"


The gross national product will therefore, be:

"\text{GNP=GDP+Net income from Broad}"

"\text{GNP}=448+3"

"\text{GNP}=451"


Income Approach.


Using Income approach the GNP is calculated as follows:

"\text{ GNP = Wages + Interest Income + Rental Income + Profit }"

"\text{GNP}=329+49+46+69"

"\text{GNP}=493"